# CAMSIC: Content-aware Masked Image Modeling Transformer for Stereo Image Compression

### v4

@@ Abstract
Existing learning-based stereo image codec adopt sophisticated transformation with simple entropy models derived from single image codecs to encode latent representations. However, those entropy models struggle to effectively capture the spatial-disparity characteristics inherent in stereo images, which leads to suboptimal rate-distortion results. In this paper, we propose a stereo image compression framework, named CAMSIC. CAMSIC independently transforms each image to latent representation and employs a powerful decoder-free Transformer entropy model to capture both spatial and disparity dependencies, by introducing a novel content-aware masked image modeling (MIM) technique. Our content-aware MIM facilitates efficient bidirectional interaction between prior information and estimated tokens, which naturally obviates the need for an extra Transformer decoder. Experiments show that our stereo image codec achieves state-of-the-art rate-distortion performance on two stereo image datasets Cityscapes and InStereo2K with fast encoding and decoding speed. Code is available at https://github.com/Xinjie-Q/CAMSIC.
@@ Introduction
$\mathcal{L}$ $=\lambda D+R$ (7)
$=\lambda\sum_{v=1}^{2}d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})+\sum_{v=1}^{2}\big{(}R(\boldsymbol{\hat{y}}_{v})+R(\boldsymbol{\hat{z}}_{v})\big{)}$
where $d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})$ represents the distortion between the original image $\boldsymbol{x}_{v}$ and reconstructed image $\boldsymbol{\hat{x}}_{v}$ under a given metric such as MSE or MS-SSIM (Wang, Simoncelli, and Bovik 2003). $R(\boldsymbol{\hat{y}}_{v})$ denotes the estimated bitrate of the quantized representation $\boldsymbol{\hat{y}}_{v}$ and $R(\boldsymbol{\hat{z}}_{v})$ denotes the estimated bitrate of the hyper latent $\boldsymbol{\hat{z}}_{v}$. $\lambda$ is a hyper parameter that is used to balance the compression rate $R$ and distortion $D$.
@@
Figure 4: Rate-distortion curves of our approach and different baselines on Cityscapes and InStereo2K datasets in PSNR and MS-SSIM.
@@
Different Entropy Models. To demonstrate the superiority of our proposed entropy model, a set of ablation experiments are conducted. Table 2 presents the BDBR results on the Cityscapes and InStereo2K datasets. Specifically, we replace our entropy model with the CNN-based ECSIC, SASIC and BCSIC entropy models as well as the content-irrelevant encoder-decoder Transformer entropy model. Note that the vanilla MIM-based Transformer entropy model outperforms the ECSIC entropy model, which is attributed to the Transformer’s inherent capability to capture long-range dependencies more effectively than CNN architectures. However, simply migrating the vanilla encoder-decoder Transformer architecture and the naive MIM technique to the entropy model leads to inefficient prior information propagation, which makes the vanilla MIM entropy model inferior to the BCSIC entropy model with bidirectional view interaction. To address this challenge and fully unleash the potential of the Transformer, we introduce a novel content-aware MIM technique to design an advanced decoder-free Transformer entropy model. The improvements demonstrate that our content-aware MIM significantly surpasses the content-irrelevant MIM, which underscores the importance of bidirectional interactions between prior information and estimated tokens in improving compression performance.
@@
Table 2: Ablation studies for different components. The first row represents our final solution and is set as the anchor to measure BD-rate. RPE means relative position embedding. ID indicates identitiy embedding.
@@ Conclusion
In this paper, we propose a stereo image compression framework, namely CAMSIC, with a succinct yet potent Transformer entropy model as the core to effectively capture the spatial-disparity correlations. Our CAMSIC consists of a straightforward encoder-decoder architecture to independently transform each image. A novel content-aware masked image modeling (MIM) technique is presented to fully exploit the power of Transformer entropy model. Our content-aware MIM unleashes bidirectional interactions between the prior information and estimated tokens, which allows us to eliminate an extra Transformer decoder. Thus, a decoder-free Transformer entropy model is proposed to improve the efficiency. Experimental results show that our approach achieves the state-of-the-art compression performance with fast encoding and decoding. Our work sheds light on the development of Transformer architecture in future image compression research.

### v5

@@ Abstract
Existing learning-based stereo image codec adopt sophisticated transformation with simple entropy models derived from single image codecs to encode latent representations. However, those entropy models struggle to effectively capture the spatial-disparity characteristics inherent in stereo images, which leads to suboptimal rate-distortion results. In this paper, we propose a stereo image compression framework, named CAMSIC. CAMSIC independently transforms each image to latent representation and employs a powerful decoder-free Transformer entropy model to capture both spatial and disparity dependencies, by introducing a novel content-aware masked image modeling (MIM) technique. Our content-aware MIM facilitates efficient bidirectional interaction between prior information and estimated tokens, which naturally obviates the need for an extra Transformer decoder. Experiments show that our stereo image codec achieves state-of-the-art rate-distortion performance on two stereo image datasets Cityscapes and InStereo2K with fast encoding and decoding speed.
@@ Introduction
Figure 4: Rate-distortion curves of our approach and different baselines on Cityscapes and InStereo2K datasets.
@@
$\mathcal{L}=\lambda\sum_{v=1}^{2}d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})+\sum_{v=1}^{2}\big{(}R(\boldsymbol{\hat{y}}_{v})+R(\boldsymbol{\hat{z}}_{v})\big{)}$ (7)
where $d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})$ represents the distortion between the original image $\boldsymbol{x}_{v}$ and reconstructed image $\boldsymbol{\hat{x}}_{v}$ under a given metric such as MSE or MS-SSIM (Wang, Simoncelli, and Bovik 2003). $R(\boldsymbol{\hat{y}}_{v})$ denotes the estimated bitrate of the quantized representation $\boldsymbol{\hat{y}}_{v}$ and $R(\boldsymbol{\hat{z}}_{v})$ denotes the estimated bitrate of the hyper latent $\boldsymbol{\hat{z}}_{v}$. $\lambda$ is a hyper parameter that is used to balance the compression rate and distortion.
@@
Different Entropy Models. Table 2 presents the BDBR results of different entropy models on the Cityscapes and InStereo2K datasets. Specifically, we replace our entropy model with the CNN-based ECSIC, SASIC and BCSIC entropy models as well as the content-irrelevant encoder-decoder Transformer entropy model. Note that the vanilla MIM-based Transformer entropy model outperforms the ECSIC entropy model, which is attributed to the Transformer’s inherent capability to capture long-range dependencies more effectively than CNN architectures. However, simply migrating the vanilla encoder-decoder Transformer architecture and the naive MIM technique to the entropy model leads to inefficient prior information propagation, which makes the vanilla MIM entropy model inferior to the BCSIC entropy model with bidirectional view interaction. To address this challenge, we introduce a novel content-aware MIM technique to design an advanced decoder-free Transformer entropy model. The improvements demonstrate that our content-aware MIM significantly surpasses the content-irrelevant MIM, which underscores the importance of bidirectional interactions between prior information and estimated tokens in improving compression performance.
@@
Table 2: Ablation studies for different components. The first row is set as the anchor to measure BD-rate. RPE means relative position embedding. ID indicates identitiy embedding.
@@ Conclusion
In this paper, we propose a stereo image compression framework, namely CAMSIC, with a succinct yet potent Transformer entropy model as the core to effectively capture the spatial-disparity correlations. Our CAMSIC consists of a straightforward encoder-decoder architecture to independently transform each image. A novel content-aware masked image modeling (MIM) technique is presented to fully exploit the power of Transformer entropy model. Our content-aware MIM unleashes bidirectional interactions between the prior information and estimated tokens, which allows us to eliminate an extra Transformer decoder. Thus, a decoder-free Transformer entropy model is proposed to improve the efficiency. Experimental results show that our approach achieves the state-of-the-art compression performance with fast encoding and decoding.